# Towards a Functionally Complete and Parameterizable TFHE Processor

### v1

@@ Abstract.
While state-of-the-art implementations focused on accelerating and outsourcing single operations, their scalability and practicality are constrained by high memory bandwidth costs. In order to overcome this, we propose an FPGA-based hardware accelerator for the evaluation of homomorphic circuits. Specifically, we design a functionally complete TFHE processor for FPGA hardware capable of processing instructions on the data completely on the FPGA. In order to achieve a higher throughput from our TFHE processor, we implement an improved programmable bootstrapping module which outperforms the current state-of-the-art by 240% to 480% more bootstrappings per second. Our efficient, compact, and scalable design lays the foundation for implementing complete FPGA-based TFHE processor architectures [^1] Open access: https://github.com/CeresB/tfhe-processor.
@@ 1. Introduction
The first successful construction of an FHE scheme was accomplished by Gentry (Gentry, 2009) in 2009. This was followed by a large body of works leveraging and improving on Gentry’s blueprint, leading to faster and more practical fully homomorphic encryption schemes. Despite these optimizations, FHE suffers from a considerable computational overhead when evaluating encrypted computations. Concretely, FHE schemes are several orders of magnitude slower than equivalent plaintext computations (Badawi et al., 2023). This severely hinders the adoption of FHE as a general solution for secure multiparty computation.
The torus-based encryption scheme TFHE (Chillotti et al., 2019) leverages the learning with errors (LWE) hardness assumption (Regev, 2005) to provide fully homomorphic encryption. TFHE relies on small parameter sets in order to provide fast homomorphic operations. Additionally, the bootstrapping operation allows the inherent encryption noise to be kept within secure and correct bounds. This is typically the most expensive operation in FHE schemes. TFHE accomplishes this by providing the fastest FHE bootstrapping algorithm, while also allowing the evaluation of arbitrary lookup tables for free during bootstrapping. This is considerable as it allows the evaluation of non-linear operations under encryption, leading to the term programmable bootstrapping (PBS).
Several works have explored hardware acceleration for FHE, particularly TFHE (Hu et al., 2024; Ye et al., 2022; Kong and Li, 2024; Van Beirendonck et al., 2023). Most focus on accelerating the PBS operation on FPGAs, offloading only this computationally intensive step while keeping other homomorphic operations on conventional hardware. Although this design enables significant PBS acceleration, it introduces high data transfer overheads to and from the FPGA, making overall performance and scalability heavily dependent on external memory bandwidth.
@@ 1.1. Related Work
Table 1. Comparison of prior work. FPT and ALT accelerate the blind rotation but do not perform a complete PBS as only the blind rotation step is performed on the FPGA (Van Beirendonck et al., 2023; Hu et al., 2024). Abbreviations: BR (Blind rotation), KS (Keyswitch)
@@
FPT (Van Beirendonck et al., 2023) FFT - - no memory computational Batched bootstrapping
ALT (Hu et al., 2024) NTT 17 & 16 - yes multiplication computational Composite NTT
KL (Kong and Li, 2024) NTT 52 yes memory bandwidth Invalid keyswitch implementation
YKP (Ye et al., 2022) NTT 33 yes multiplication bandwidth -
MXP (Gener et al., 2021) FFT - N/A N/A N/A Small-scale CPU-like architecture with special instructions to accelerate common TFHE operations
Ohba (Ohba et al., 2025) NTT - N/A N/A N/A CPU-like architecture with special instruction to accelerate NTT computation
@@
Gener et al. (Gener et al., 2021) propose an accelerated general-purpose vector engine and demonstrate its capabilities by accelerating the polynomial multiplication used for TFHE bootstrapping. Their solution is not specific to TFHE and performs considerably slower than comparable CPU implementations. Ohba et al. (Ohba et al., 2025) similarly present a general-purpose NTT acceleration engine, which is then applied to TFHE. Similarly, Turbo-FHE (Nassar et al., 2025) leverages the Karatsuba algorithm to accelerate polynomial multiplication on FPGAs. The authors do not provide sufficient details for a comparison, nor do they provide comparisons themselves. FPT (Van Beirendonck et al., 2023) accelerates polynomial multiplication by using fixed-point arithmetic for the Fast Fourier Transform (FFT). However, this approach limits numerical precision, increasing the probability of decryption failures. Similar to our design, FPT leverages on-chip High-Bandwidth Memory (HBM) to mitigate external memory bandwidth constraints. Furthermore, its batched operation scheduling provides additional time between memory accesses, effectively reducing overall bandwidth demand.
ALT (Hu et al., 2024), Kong and Li (KL) (Kong and Li, 2024), as well as Ye et al. (YKP) (Ye et al., 2022), accelerate the programmable bootstrapping operation by leveraging key-unrolling and performing several small NTTs in parallel. Additionally, ALT uses two small primes to perform composite NTTs, leveraging the Chinese remainder theorem. All three works rely on external off-chip memory to store the bootstrapping key, adding large memory bandwidth constraints. ALT solves this problem by prefetching parts of the key during computation. The key-switching operation implemented by Kong and Li requires the secret key to be present in the clear on the FPGA, compromising the security of the implementation.
FHE Processors beyond TFHE: Several promising efforts have explored dedicated FHE processor designs for schemes beyond TFHE. Specifically, BASALISC (Geelen et al., 2023) targets the BGV scheme, while F1 (Samardzic et al., 2021) introduces a general-purpose FHE processor compatible with BGV, CKKS, and general polynomial operations (Nabeel et al., 2024). These efforts present ASIC-based designs emphasizing efficient memory management, hardware reuse for compactness, and support for diverse FHE parameter sets. Each design also defines its own scheme-specific FHE instruction set. However, these designs remain limited in their applicability to TFHE, as their architectures and instruction sets are optimized for FHE schemes which function fundamentally different to TFHE regarding their parameters and capabilities.
@@ 2.2. Torus-based Fully homomorphic encryption
Fully homomorphic encryption (FHE) schemes enable the evaluation of arbitrary functions on encrypted data. This is typically accomplished by constructing schemes that allow both unlimited additions and multiplications under encryption, thus providing functional completeness. All major FHE schemes rely on noise for encryption. This noise grows with each homomorphic operation, reducing the probability of correct decryption. The introduction of the bootstrapping operation to manage this noise led to the construction of the first FHE scheme by Gentry (Gentry, 2009).
The vast majority of FHE schemes rely on the learning with errors (LWE) hardness assumption (Regev, 2005), and its ring-variant RLWE (Lyubashevsky et al., 2013), for encryption. The generalized LWE (GLWE) problems capture both variants as illustrated by Definition 2.1.
@@
(GLWE (Regev, 2005; Lyubashevsky et al., 2013)) Let $k\geq 1$ be an integer and $N$ a power of two and $X^{N}+1$ the corresponding cyclotomic polynomial. Let $\chi$ be a Gaussian distribution over $\mathbb{T}_{N}[X]$ with coefficient-wise variance $\sigma^{2}$, and let $s\in\mathbb{B}_{N}[X]$. Let $\text{GLWE}_{s,\chi}$ denote the distribution over $\mathbb{T}_{N}[X]^{k}\times\mathbb{T}_{N}[X]$ obtained by sampling a uniform vector $\mathbf{a}\overset{\mathdollar}{\leftarrow}\mathbb{T}_{N}[X]^{k}$ and yielding the tuple $(\mathbf{a},b)$ with $b=\langle\mathbf{a},\mathbf{s}\rangle+e$, where $e\leftarrow\chi$.
@@
The torus-based FHE scheme TFHE (Chillotti et al., 2019) encrypts messages by encoding them into the noise of an GLWE sample. A message $m\in\mathfrak{R}_{p}^{N}$ is encrypted under the secret key $\texttt{sk}\in\mathbb{B}_{N}[X]^{k}$ as follows:
@@
TFHE (Chillotti et al., 2019) leverages LWE ciphertexts to encrypt integers. It provides addition, multiplication, as well as programmable bootstrapping (PBS). This differs from traditional bootstrapping by allowing the evaluation of an arbitrary lookup table for free during bootstrapping. RLWE ciphertexts are only applied during the computation of the PBS operation.
@@ 2.3. Number Theoretic Transform
The discrete Fourier transform (DFT) (Satriawan et al., 2023) enables efficient polynomial multiplication. Polynomials represented in the Fourier domain can be multiplied by simple element-wise multiplication, as opposed to the convolution of coefficients otherwise required. Let $g$ be a polynomial with coefficients $g_{0},\dots,g_{N-1}$. An $N$-point DFT is defined as the evaluation of $g$ at the $N$-th roots of unity generated by the primitive $N$-th root of unity $\omega$. That is
@@ 2.4. Field Programmable Gate Arrays
Field-Programmable Gate Arrays (FPGAs) (Kuon and Rose, 2009) are programmable integrated circuits composed of arithmetic and logic units, memory blocks, and routing interconnects. They offer high computational performance and energy efficiency compared to general-purpose hardware such as CPUs and GPUs, while remaining far more flexible and cost-effective than ASICs, whose development is expensive and time-consuming (Jayasena and Mishra, 2024).
FPGAs are also well-suited for hardware prototyping and testing, as their reprogrammable nature allows rapid iteration and validation of hardware designs before committing to costly ASIC fabrication (Jayasena and Mishra, 2024). An FPGA consists mainly of digital signal processing (DSP) units for arithmetic, lookup tables (LUTs) for logic, flip-flops (FFs) for bit storage, and various on-chip RAMs. Designs are written in a hardware description language (HDL) and synthesized into routing configurations that define signal paths between elements. FPGA performance depends on the operating frequency, which is constrained by signal propagation delays across routed connections. Therefore, efficient design requires minimizing long interconnects and balancing resource utilization to achieve high clock speeds and reliable operation.
@@ 3.1. Modular Reduction
Efficient Reduction Algorithms: A naive reduction by division is too costly for hardware implementation. The Barrett reduction (Barrett, 1987) offers an efficient alternative for arbitrary moduli and can be combined with Montgomery multiplication for modular products. However, both methods still require complex arithmetic and are less suited to high-throughput hardware. Instead, we leverage moduli that simplify reduction. Power-of-two moduli allow implicit reduction via overflow, while Mersenne and Solinas primes enable reduction using only shifts and additions. Such moduli significantly reduce hardware complexity and latency.
@@ 4. Towards a TFHE processor
Note that our design does not include a general external product for ciphertext-ciphertext multiplication. Instead, we assume that homomorphic multiplications are performed by concatenating messages and performing a programmable bootstrap with an accordingly constructed lookup table. This is also the approach followed by the tfhe-rs library (Zama, 2022b).
@@ 4.1. External Bandwidth Requirements
Assuming a 64-bit address space, a processor instruction operation requires $3\cdot 64$ bits for the data addresses and $\log N$ bits specifically required for sample extraction. Finally, additional bits are required for the instructions themselves. Two bits suffice for this, assuming the minimal instruction set PBS, AddMul, and KS. Note that the keys are stored on the FPGA, saving additional external bandwidth requirements. Considering the largest parameter sets tested, as well as the highest rate of PBSs per second achieved by our design, this results in an external bandwidth requirement of around 3 Megabits per second. Prior work requires several Gigabytes per second (Hu et al., 2024; Ye et al., 2022; Kong and Li, 2024).
@@ 4.2. Internal Bandwidth Requirements
The highest internal bandwidth requirement incurred by our design is from the bootstrapping key BSK. During bootstrapping, different parts of the key are required. The underlying RGSW ciphertexts are quite considerable in size. Loading of the bootstrapping key is amortized through batched bootstrapping as proposed in FPT (Van Beirendonck et al., 2023). This keeps a key element longer in use, allowing more clock cycles for the subsequent element to be loaded from memory. Setting the batch size trades lower memory consumption and latency for a higher internal memory bandwidth requirement.
@@ 5.1. Experimental Setup
Table 2. Resources of the FPGAs used by different implementations. All FPGAs additionally feature multiple GB off-chip DDR4 RAM. SRAM is BRAM + URAM + PLRAM. RAM unit conversions: 1 LUTRAM = 64 bit, 1 BRAM = 36 kb, 1 URAM = 288 kb. *LUTRAM for the FPGA used by YKP is estimated based on the fact that one half of CLB LUTs are LUTRAM (AMD, 2024b) and the VU13P has 1728 K CLB LUTs (AMD, 2024a). Abbreviations: DCAC (Data Center Accelerator Card), VU+ (Virtex UltraScale+ HBM)
@@
FPT (Van Beirendonck et al., 2023) AMD Alveo U280 DCAC 1304 K 2607 K 9024 N/A 2016 960 24 Mb 365 Mb 8 GB HBM with up to 460 GB/s
ALT (Hu et al., 2024) AMD Alveo U250 DCAC 1728 K 3456 K 12288 N/A 2688 1280 N/A 452 Mb External memory with up to 77 GB/s
KL (Kong and Li, 2024) AMD VU+ VCU128 1304 K 2607 K 9024 601 K 2016 960 0 Mb 341 Mb 8 GB HBM with up to 460 GB/s
YKP (Ye et al., 2022) AMD VU+ VU13P 1728 K 3456 K 12288 864 K* 2688 1280 0 Mb 452 Mb External memory with up to 77 GB/s
Ohba (Ohba et al., 2025) AMD XCVU47P 1304 K 2607 K 9024 601 K 2016 960 0 Mb 341 Mb 16 GB HBM with up to 460 GB/s

### v2

@@ Abstract.
While state-of-the-art implementations focused on accelerating and outsourcing single operations, their scalability and practicality are constrained by high memory bandwidth costs. In order to overcome this, we propose an FPGA-based hardware accelerator for the evaluation of homomorphic circuits. Specifically, we design a functionally complete TFHE processor for FPGA hardware capable of processing instructions on the data completely on the FPGA. In order to achieve a higher throughput from our TFHE processor, we implement an improved programmable bootstrapping module which outperforms the current state-of-the-art by 240% to 480% more bootstrappings per second. Our efficient, compact, and scalable design lays the foundation for implementing complete FPGA-based TFHE processor architectures [^1] Open access: https://github.com/CeresB/tfhe-processor-artifacts.
@@ 1. Introduction
The first successful construction of an FHE scheme was accomplished by Gentry (gentry09) in 2009. This was followed by a large body of works leveraging and improving on Gentry’s blueprint, leading to faster and more practical fully homomorphic encryption schemes. Despite these optimizations, FHE suffers from a considerable computational overhead when evaluating encrypted computations. Concretely, FHE schemes are several orders of magnitude slower than equivalent plaintext computations (entry_whitepaper). This severely hinders the adoption of FHE as a general solution for secure multiparty computation.
The torus-based encryption scheme TFHE (TFHE_original_paper) leverages the learning with errors (LWE) hardness assumption (Regev2005) to provide fully homomorphic encryption. TFHE relies on small parameter sets in order to provide fast homomorphic operations. Additionally, the bootstrapping operation allows the inherent encryption noise to be kept within secure and correct bounds. This is typically the most expensive operation in FHE schemes. TFHE accomplishes this by providing the fastest FHE bootstrapping algorithm, while also allowing the evaluation of arbitrary lookup tables for free during bootstrapping. This is considerable as it allows the evaluation of non-linear operations under encryption, leading to the term programmable bootstrapping (PBS).
Several works have explored hardware acceleration for FHE, particularly TFHE (ALT; Tian_Ye; Tianqi_Kong; FPT). Most focus on accelerating the PBS operation on FPGAs, offloading only this computationally intensive step while keeping other homomorphic operations on conventional hardware. Although this design enables significant PBS acceleration, it introduces high data transfer overheads to and from the FPGA, making overall performance and scalability heavily dependent on external memory bandwidth.
@@ 1.1. Related Work
Table 1. Comparison of prior work. FPT and ALT accelerate the blind rotation but do not perform a complete PBS as only the blind rotation step is performed on the FPGA (FPT; ALT). Abbreviations: BR (Blind rotation), KS (Keyswitch)
@@
FPT (FPT) FFT - - no memory computational Batched bootstrapping
ALT (ALT) NTT 17 & 16 - yes multiplication computational Composite NTT
KL (Tianqi_Kong) NTT 52 yes memory bandwidth Invalid keyswitch implementation
YKP (Tian_Ye) NTT 33 yes multiplication bandwidth -
MXP (MXP) FFT - N/A N/A N/A Small-scale CPU-like architecture with special instructions to accelerate common TFHE operations
Ohba (Ohba2025_nvme) NTT - N/A N/A N/A CPU-like architecture with special instruction to accelerate NTT computation
@@
Gener et al. (MXP) propose an accelerated general-purpose vector engine and demonstrate its capabilities by accelerating the polynomial multiplication used for TFHE bootstrapping. Their solution is not specific to TFHE and performs considerably slower than comparable CPU implementations. Ohba et al. (Ohba2025_nvme) similarly present a general-purpose NTT acceleration engine, which is then applied to TFHE. Similarly, Turbo-FHE (Turbo_FHE) leverages the Karatsuba algorithm to accelerate polynomial multiplication on FPGAs. The authors do not provide sufficient details for a comparison, nor do they provide comparisons themselves. FPT (FPT) accelerates polynomial multiplication by using fixed-point arithmetic for the Fast Fourier Transform (FFT). However, this approach limits numerical precision, increasing the probability of decryption failures. Similar to our design, FPT leverages on-chip High-Bandwidth Memory (HBM) to mitigate external memory bandwidth constraints. Furthermore, its batched operation scheduling provides additional time between memory accesses, effectively reducing overall bandwidth demand.
ALT (ALT), Kong and Li (KL) (Tianqi_Kong), as well as Ye et al. (YKP) (Tian_Ye), accelerate the programmable bootstrapping operation by leveraging key-unrolling and performing several small NTTs in parallel. Additionally, ALT uses two small primes to perform composite NTTs, leveraging the Chinese remainder theorem. All three works rely on external off-chip memory to store the bootstrapping key, adding large memory bandwidth constraints. ALT solves this problem by prefetching parts of the key during computation. The key-switching operation implemented by Kong and Li requires the secret key to be present in the clear on the FPGA, compromising the security of the implementation.
FHE Processors beyond TFHE: Several promising efforts have explored dedicated FHE processor designs for schemes beyond TFHE. Specifically, BASALISC (BASALISC) targets the BGV scheme, while F1 (F1) introduces a general-purpose FHE processor compatible with BGV, CKKS, and general polynomial operations (CoFHEE_extended). These efforts present ASIC-based designs emphasizing efficient memory management, hardware reuse for compactness, and support for diverse FHE parameter sets. Each design also defines its own scheme-specific FHE instruction set. However, these designs remain limited in their applicability to TFHE, as their architectures and instruction sets are optimized for FHE schemes which function fundamentally different to TFHE regarding their parameters and capabilities.
@@ 2.2. Torus-based Fully homomorphic encryption
Fully homomorphic encryption (FHE) schemes enable the evaluation of arbitrary functions on encrypted data. This is typically accomplished by constructing schemes that allow both unlimited additions and multiplications under encryption, thus providing functional completeness. All major FHE schemes rely on noise for encryption. This noise grows with each homomorphic operation, reducing the probability of correct decryption. The introduction of the bootstrapping operation to manage this noise led to the construction of the first FHE scheme by Gentry (gentry09).
The vast majority of FHE schemes rely on the learning with errors (LWE) hardness assumption (Regev2005), and its ring-variant RLWE (lyubashevsky_ideal_2012), for encryption. The generalized LWE (GLWE) problems capture both variants as illustrated by Definition 2.1.
@@
(GLWE (Regev2005; lyubashevsky_ideal_2012)) Let $k\geq 1$ be an integer and $N$ a power of two and $X^{N}+1$ the corresponding cyclotomic polynomial. Let $\chi$ be a Gaussian distribution over $\mathbb{T}_{N}[X]$ with coefficient-wise variance $\sigma^{2}$, and let $s\in\mathbb{B}_{N}[X]$. Let $\text{GLWE}_{s,\chi}$ denote the distribution over $\mathbb{T}_{N}[X]^{k}\times\mathbb{T}_{N}[X]$ obtained by sampling a uniform vector $\mathbf{a}\overset{\mathdollar}{\leftarrow}\mathbb{T}_{N}[X]^{k}$ and yielding the tuple $(\mathbf{a},b)$ with $b=\langle\mathbf{a},\mathbf{s}\rangle+e$, where $e\leftarrow\chi$.
@@
The torus-based FHE scheme TFHE (TFHE_original_paper) encrypts messages by encoding them into the noise of an GLWE sample. A message $m\in\mathfrak{R}_{p}^{N}$ is encrypted under the secret key $\texttt{sk}\in\mathbb{B}_{N}[X]^{k}$ as follows:
@@
TFHE (TFHE_original_paper) leverages LWE ciphertexts to encrypt integers. It provides addition, multiplication, as well as programmable bootstrapping (PBS). This differs from traditional bootstrapping by allowing the evaluation of an arbitrary lookup table for free during bootstrapping. RLWE ciphertexts are only applied during the computation of the PBS operation.
@@ 2.3. Number Theoretic Transform
The discrete Fourier transform (DFT) (ntt_guide) enables efficient polynomial multiplication. Polynomials represented in the Fourier domain can be multiplied by simple element-wise multiplication, as opposed to the convolution of coefficients otherwise required. Let $g$ be a polynomial with coefficients $g_{0},\dots,g_{N-1}$. An $N$-point DFT is defined as the evaluation of $g$ at the $N$-th roots of unity generated by the primitive $N$-th root of unity $\omega$. That is
@@ 2.4. Field Programmable Gate Arrays
Field-Programmable Gate Arrays (FPGAs) (fpga_vs_asic_basics) are programmable integrated circuits composed of arithmetic and logic units, memory blocks, and routing interconnects. They offer high computational performance and energy efficiency compared to general-purpose hardware such as CPUs and GPUs, while remaining far more flexible and cost-effective than ASICs, whose development is expensive and time-consuming (jayasena2024directed).
FPGAs are also well-suited for hardware prototyping and testing, as their reprogrammable nature allows rapid iteration and validation of hardware designs before committing to costly ASIC fabrication (jayasena2024directed). An FPGA consists mainly of digital signal processing (DSP) units for arithmetic, lookup tables (LUTs) for logic, flip-flops (FFs) for bit storage, and various on-chip RAMs. Designs are written in a hardware description language (HDL) and synthesized into routing configurations that define signal paths between elements. FPGA performance depends on the operating frequency, which is constrained by signal propagation delays across routed connections. Therefore, efficient design requires minimizing long interconnects and balancing resource utilization to achieve high clock speeds and reliable operation.
@@ 3.1. Modular Reduction
Efficient Reduction Algorithms: A naive reduction by division is too costly for hardware implementation. The Barrett reduction (barrett) offers an efficient alternative for arbitrary moduli and can be combined with Montgomery multiplication for modular products. However, both methods still require complex arithmetic and are less suited to high-throughput hardware. Instead, we leverage moduli that simplify reduction. Power-of-two moduli allow implicit reduction via overflow, while Mersenne and Solinas primes enable reduction using only shifts and additions. Such moduli significantly reduce hardware complexity and latency.
@@ 4. Towards a TFHE processor
Note that our design does not include a general external product for ciphertext-ciphertext multiplication. Instead, we assume that homomorphic multiplications are performed by concatenating messages and performing a programmable bootstrap with an accordingly constructed lookup table. This is also the approach followed by the tfhe-rs library (tfhe_rs).
@@ 4.1. External Bandwidth Requirements
Assuming a 64-bit address space, a processor instruction operation requires $3\cdot 64$ bits for the data addresses and $\log N$ bits specifically required for sample extraction. Finally, additional bits are required for the instructions themselves. Two bits suffice for this, assuming the minimal instruction set PBS, AddMul, and KS. Note that the keys are stored on the FPGA, saving additional external bandwidth requirements. Considering the largest parameter sets tested, as well as the highest rate of PBSs per second achieved by our design, this results in an external bandwidth requirement of around 3 Megabits per second. Prior work requires several Gigabytes per second (ALT; Tian_Ye; Tianqi_Kong).
@@ 4.2. Internal Bandwidth Requirements
The highest internal bandwidth requirement incurred by our design is from the bootstrapping key BSK. During bootstrapping, different parts of the key are required. The underlying RGSW ciphertexts are quite considerable in size. Loading of the bootstrapping key is amortized through batched bootstrapping as proposed in FPT (FPT). This keeps a key element longer in use, allowing more clock cycles for the subsequent element to be loaded from memory. Setting the batch size trades lower memory consumption and latency for a higher internal memory bandwidth requirement.
@@ 5.1. Experimental Setup
Table 2. Resources of the FPGAs used by different implementations. All FPGAs additionally feature multiple GB off-chip DDR4 RAM. SRAM is BRAM + URAM + PLRAM. RAM unit conversions: 1 LUTRAM = 64 bit, 1 BRAM = 36 kb, 1 URAM = 288 kb. *LUTRAM for the FPGA used by YKP is estimated based on the fact that one half of CLB LUTs are LUTRAM (vivado_documentation_lutram) and the VU13P has 1728 K CLB LUTs (vu13_datasheet). Abbreviations: DCAC (Data Center Accelerator Card), VU+ (Virtex UltraScale+ HBM)
@@
FPT (FPT) AMD Alveo U280 DCAC 1304 K 2607 K 9024 N/A 2016 960 24 Mb 365 Mb 8 GB HBM with up to 460 GB/s
ALT (ALT) AMD Alveo U250 DCAC 1728 K 3456 K 12288 N/A 2688 1280 N/A 452 Mb External memory with up to 77 GB/s
KL (Tianqi_Kong) AMD VU+ VCU128 1304 K 2607 K 9024 601 K 2016 960 0 Mb 341 Mb 8 GB HBM with up to 460 GB/s
YKP (Tian_Ye) AMD VU+ VU13P 1728 K 3456 K 12288 864 K* 2688 1280 0 Mb 452 Mb External memory with up to 77 GB/s
Ohba (Ohba2025_nvme) AMD XCVU47P 1304 K 2607 K 9024 601 K 2016 960 0 Mb 341 Mb 16 GB HBM with up to 460 GB/s